OpenStudy (anonymous):

how do you factor a^2^n +6a^n+8

OpenStudy (abb0t):

Just to be clear, is it really a\(\sf\color{red}{^{2n}}\)+a\(\sf\color{red}{^{n}}\)+8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6a^n sorry

OpenStudy (abb0t):

Rewrite it using LaTex (the equation editor) on the bottom of this little box.

OpenStudy (anonymous):

\[a ^{2n}+6a ^{n}+8\]

OpenStudy (abb0t):

Well, are you solving for \(a\) or for \(n\)?

OpenStudy (anonymous):

It says to factor completely or indicate if prime

OpenStudy (abb0t):

Well, eitheway, what two numbers when multiplied give you 8, but when added equal 6?

OpenStudy (abb0t):

I am going to assume that n = some number (for simplicity, i will assume n = 1)

OpenStudy (anonymous):

ok

OpenStudy (abb0t):

Are you with me so far? B/c 6 or 8 is nOT the answer.

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

Well, what are those two numbers @red32968 ___ \(\times\) ___ = 8 ___ + ___ = 6

OpenStudy (anonymous):

4*2 4+2

OpenStudy (abb0t):

Yes, Correct! So you have: a\(^{2n}\)+4a\(^n\)+2a\(^n\)+8, next you will group them (a\(^{2n}\)+4a\(^n\))+(2a\(^n\)+8). Notice you can factor out a few terms for them. Correct? Factor those terms out for me.

OpenStudy (abb0t):

HINT: x\(^{2b}\)+x\(^b\) = x\(^b\)(x\(^b\)+1)....it's good to remember your exponent rules here.

OpenStudy (abb0t):

NOTE: x\(^b\)(x\(^b\))= x\(^{b+b}\)= x\(^{2b}\)

OpenStudy (anonymous):

\[a ^{2n}(6^{an}+8)\]a ^{2n+n}

OpenStudy (abb0t):

No, not quite.

OpenStudy (anonymous):

where did I make the mistake

OpenStudy (anonymous):

\[6a ^{n}\]

OpenStudy (abb0t):

Here's another HINT: a\(^n\)(a\(^n\)+4) that is ONE of them.

OpenStudy (anonymous):

the second should be \[a ^{n}(a ^{n}+8)\]

OpenStudy (abb0t):

No! You can ONLY factor out a 2 from the second one!

OpenStudy (anonymous):

2\[2(a ^{n}+8)\]

OpenStudy (anonymous):

actually IT SHOULD BE +4

OpenStudy (abb0t):

YES! Correct!

OpenStudy (abb0t):

So, now that you have everything factored, what is your final answer? I can compare it.

OpenStudy (abb0t):

(__ + __)(__ +__) what are the __ ?

OpenStudy (anonymous):

\[A ^{N}(A ^{n}+4) 2 ( a^{n}+4)\]

OpenStudy (abb0t):

Not quite, you forgot the +2, but the final answer is: \(\sf\color{red}{(a^n+2)(a^n+4)}\)

OpenStudy (anonymous):

thank you for your patience and help